# Energy-based generator matching: A neural sampler for general state space

### v2

@@ 4 Experiments
Algorithm $\downarrow \mathcal{E}$-$\mathcal{W}_{1}$downarrow $M$-$\mathcal{W}_{1}$downarrow $\mathcal{E}$-$\mathcal{W}_{1}$downarrow $M$-$\mathcal{W}_{1}$downarrow $\mathcal{E}$-$\mathcal{W}_{1}$downarrow $M$-$\mathcal{W}_{1}$downarrow $\mathcal{E}$-$\mathcal{W}_{1}$downarrow $M$-$\mathcal{W}_{1}$downarrow
Gibbs 0.10±$\pm 0.02$ 0.06±$\pm 0.02$ 0.71±$\pm 0.43$ 0.23±$\pm 0.18$ 0.53±$\pm 0.04$ 0.04±$\pm 0.01$ 5.29±$\pm 1.95$ 0.29±$\pm 0.11$
EGM (ours) 0.20±$\pm 0.06$ 0.02±$\pm 0.01$ 3.73±$\pm 0.39$ 0.24±$\pm 0.02$ 0.84±$\pm 0.07$ 0.02±$\pm 0.00$ 19.94±$\pm 0.69$ 0.36±$\pm 0.01$
+Bootstrapping (ours) 0.10±$\pm 0.06$ 0.02±$\pm 0.01$ 0.60±$\pm 0.12$ 0.04±$\pm 0.01$ 0.39±$\pm 0.34$ 0.06±$\pm 0.07$ 2.51±$\pm 0.16$ 0.24±$\pm 0.01$
@@
Our primary evaluation metric is the energy-$\mathcal{W}_{1}$ ($\mathcal{E}$-$\mathcal{W}_{1}$) distance, consistently applied across tasks and complemented by qualitative visualizations. Since no jump-based or hybrid discrete–continuous neural sampler baselines exist, we include a traditional Gibbs sampler [16] as our competitive baseline. Although Gibbs sampling is guaranteed to converge given sufficiently many iterations, we use four parallel chains with 6000 steps each to reflect a realistic computational budget for the per-sample cost and to yield reasonably competitive results. Detailed descriptions of the energy functions, evaluation metrics, and experimental protocols are provided in Appendix D.
4.1 Discrete EGM on the 2D Ising Model
@@
We report the quantitative results in Table 2, alongside comparisons to the traditional parallel-Gibbs sampler as a baseline. Given the absence of metric structure in the discrete Ising model, we adopt the $\mathcal{W}_{1}$ distance over energy and average magnetization $M=\sum_{i}x_{i}$ as metrics. EGM consistently matches or exceeds baseline performance across metrics. Especially, bootstrapping improves the performance even in high-dimensional settings near critical temperatures $\beta=0.4$.
Qualitative evaluations presented in Figure 2 depict energy and average magnetization histograms for the Ising model with low temperature $\beta=0.4$, showing EGM’s ability to accurately capture the true energy distribution.
@@ 4.2 Multimodal EGM on GB-RBM, JointDW4 and JointMoG
Algorithm $\downarrow \mathcal{E}$-$\mathcal{W}_{1}$downarrow $x$-$\mathcal{W}_{2}$downarrow $\mathcal{E}$-$\mathcal{W}_{1}$downarrow $\mathcal{E}$-$\mathcal{W}_{1}$downarrow
Gibbs 0.77±$\pm 0.03$ 5.12±$\pm 0.21$ 4.80±$\pm 0.02$ 7.76±$\pm 0.02$
EGM (ours) 0.33±$\pm 0.02$ 0.63±$\pm 0.10$ 2.45±$\pm 0.07$ 8.83±$\pm 0.12$
+Bootstrapping (ours) 0.20±$\pm 0.15$ 0.61±$\pm 0.09$ 1.65±$\pm 0.96$ 1.20±$\pm 0.32$
@@
We validate EGM on joint discrete–continuous sampling tasks using three synthetic benchmarks: Gaussian-Bernoulli restricted Boltzmann machine (GB-RBM) [9], double-well potential with type-dependent interactions (JointDW4), and discrete-continuous joint mixture of Gaussians (JointMoG). We choose GB-RBM to show an application of sampling from the energy-based model. JointDW4 mimics a simplified molecular sequence-structure co-generation problem. JointMoG is a mixed-state extension of MoG that is common in diffusion-sampler benchmarks. For all the experiments, we adopt conditional-OT or variance exploding paths as a conditional probability path for the continuous flow sampler, and a masked diffusion path for the discrete jump sampler.
@@
Quantitative outcomes are summarized in Table 3. For GB-RBM, we also use $x$-$\mathcal{W}_{2}$ for the first two continuous dimensions. Qualitative results for GB-RBM and JointMoG in Figure 3 visually demonstrate EGM’s capacity to capture all distinct modes accurately. In Figure 4, we observe that EGM (with bootstrapping) correctly models the true energy distribution in JointMoG compared to the Gibbs sampler. Bootstrapping consistently improves performance in capturing multiple modes of the distributions.
@@ 5 Related Works
Diffusion samplers. Diffusion samplers cast sampling as a stochastic control or denoising problem. Early work, such as the path integral sampler (PIS) [41], required on-policy SDE simulations per update, leading to high computational costs. Subsequent variants explored different probability paths [37, 5, 38] and divergences [38, 31], yet most still rely on solving or simulating SDEs to compute their objectives. To reduce this burden, off-policy methods like iDEM [1], iEFM [39], and BNEM [30] estimate the marginal score (or flow) directly and eliminate inner-loop simulation; PINN-based approaches such as NETS [2] similarly train samplers without rollouts. While these advances improve efficiency in continuous spaces, discrete or jump process samplers remain underexplored. Recently, LEAPS [20] introduced a proactive importance sampling scheme for discrete state spaces using locally equivariant networks. However, it requires many energy evaluations at inference since it uses escorted transport.
Generative flow networks (GFNs). GFNs [3, 22] formulate sampling as a sequential decision process: a forward policy constructs an object step by step, and a backward policy enforces consistency so that trajectories terminate with probability proportional to a target reward (or unnormalized density). Training objectives like trajectory balance [26] and detailed balance [4] enable off-policy updates. Furthermore, Falet et al. [12] propose the sampler for a sparse graphical model with a local objective similar to our forward-looking parametrization. Though diffusion sampler is a special case of a continuous-state continuous-time GFNs [40, 6], it remains an interesting question that training a CTMP on a general state space, as ours, can be unified under the GFNs framework.
Relation to iDEM and BNEM. In EGM, choosing $q_{1|t}(x_{1}|x)\propto p_{t|1}(x|x_{1})$ yields simple weights $\tilde{w}(x,x_{1})=\tilde{p}_{1}(x_{1})=\exp(-\mathcal{E}(x_{1}))$. In the VE path on continuous space where target score identity [11] is available, the integrand can be replaced with the target score instead of the conditional score. This yields the same estimator used in iDEM [1] and NEM [30]. Yet, it remains an open question whether it’s possible to derive the target score identity on a general state space. Recently, Zhang et al. [42] introduce the target score identity on discrete states with neighborhood structure, which might improve the accuracy of our estimator on discrete space when combined.
@@ A.1 Importance sampling for the generator estimation
This section provides a detailed derivation of the importance sampling estimator for the generator and its parametrization presented in Equation 9.
Existence of densities. Consider the measurable space $(S,\Sigma)$ with a $\Sigma$-measurable reference measure $\nu$ (e.g., counting measure if $S$ is discrete and Lebesgue measure if $S=\mathbb{R}^{d}$), as introduced in the Section 2. Since we focus on a sampling problem, the target distribution $p_{1}$ admits a $\nu$-density $\frac{dp_{1}}{d\nu}:S\to\mathbb{R}_{\geq 0}$. Assume that the joint probability measure $p_{t,1}$ is absolutely continuous with respect to the product measure $\nu\otimes\nu$. Then, all related probability measures $p_{t},p_{t|1},p_{1|t}$ admits corresponding $\nu$-densities, expressed as:
@@
SNIS estimation of the generator. We introduce a proposal distribution $q_{1|t}:\Sigma\times S\to\mathbb{R}_{\geq 0}$, satisfying absolute continuity conditions: $p_{1|t}(\cdot|x)\ll q_{1|t}(\cdot|x)$, $q_{1|t}(\cdot|x)\ll\nu$ and $\nu\ll q_{1|t}(\cdot|x)$ for all $x\in S$. Using the marginalization trick Equation 3 and the Radon-Nikodym theorem [23, Chapter 7], we have:
@@ A.2 Proof of Theorem 1
Theorem 2 (Restatement of Theorem 1)
@@
The KFE is satisfied for the conditional probability path $p{t|r}$ by definition:
@@
where the marginal consistency assumption Equation 13 is applied in the third equality. This concludes the proof.
@@ A.3 Derivation of bootstrapping estimator for generator estimation
We derive a bootstrapping estimator for the marginal generator and its parametrization proposed in Equation 21, based on the marginalization trick in Equation 54.

### v3

@@ 4 Experiments
Our primary evaluation metric is the energy-$\mathcal{W}_{1}$ ($\mathcal{E}$-$\mathcal{W}_{1}$) distance, consistently applied across tasks and complemented by qualitative visualizations. Since no jump-based or hybrid discrete–continuous neural sampler baselines exist, we include a traditional Gibbs sampler [geman1984stochastic] as our competitive baseline. Although Gibbs sampling is guaranteed to converge given sufficiently many iterations, we use four parallel chains with 6000 steps each to reflect a realistic computational budget for the per-sample cost and to yield reasonably competitive results. Detailed descriptions of the energy functions, evaluation metrics, and experimental protocols are provided in Appendix˜D.
4.1 Discrete EGM on the 2D Ising Model
@@
Algorithm $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$ $M$-$\mathcal{W}_{1}$ $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$ $M$-$\mathcal{W}_{1}$ $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$ $M$-$\mathcal{W}_{1}$ $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$ $M$-$\mathcal{W}_{1}$ $\downarrow$
[…]
@@
We report the quantitative results in Table˜2, alongside comparisons to the traditional parallel-Gibbs sampler as a baseline. Given the absence of metric structure in the discrete Ising model, we adopt the $\mathcal{W}_{1}$ distance over energy and average magnetization $M=\sum_{i}x_{i}$ as metrics. EGM consistently matches or exceeds baseline performance across metrics. Especially, bootstrapping improves the performance even in high-dimensional settings near critical temperatures $\beta=0.4$.
Qualitative evaluations presented in Figure˜2 depict energy and average magnetization histograms for the Ising model with low temperature $\beta=0.4$, showing EGM’s ability to accurately capture the true energy distribution.
@@ 4.2 Multimodal EGM on GB-RBM, JointDW4 and JointMoG
Algorithm $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$ $x$-$\mathcal{W}_{2}$ $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$ $\mathcal{E}$-$\mathcal{W}_{1}$ $\downarrow$
[…]
@@
We validate EGM on joint discrete–continuous sampling tasks using three synthetic benchmarks: Gaussian-Bernoulli restricted Boltzmann machine (GB-RBM) [cho2011improved], double-well potential with type-dependent interactions (JointDW4), and discrete-continuous joint mixture of Gaussians (JointMoG). We choose GB-RBM to show an application of sampling from the energy-based model. JointDW4 mimics a simplified molecular sequence-structure co-generation problem. JointMoG is a mixed-state extension of MoG that is common in diffusion-sampler benchmarks. For all the experiments, we adopt conditional-OT or variance exploding paths as a conditional probability path for the continuous flow sampler, and a masked diffusion path for the discrete jump sampler.
@@
Quantitative outcomes are summarized in Table˜3. For GB-RBM, we also use $x$-$\mathcal{W}_{2}$ for the first two continuous dimensions. Qualitative results for GB-RBM and JointMoG in Figure˜3 visually demonstrate EGM’s capacity to capture all distinct modes accurately. In Figure˜4, we observe that EGM (with bootstrapping) correctly models the true energy distribution in JointMoG compared to the Gibbs sampler. Bootstrapping consistently improves performance in capturing multiple modes of the distributions.
@@ 5 Related Works
Diffusion samplers. Diffusion samplers cast sampling as a stochastic control or denoising problem. Early work, such as the path integral sampler (PIS) [zhang2021path], required on-policy SDE simulations per update, leading to high computational costs. Subsequent variants explored different probability paths [vargas2023denoising, berner2022optimal, vargas2023transport] and divergences [vargas2023transport, richter2023improved], yet most still rely on solving or simulating SDEs to compute their objectives. To reduce this burden, off-policy methods like iDEM [akhound2024iterated], iEFM [woo2024iterated], and BNEM [ouyang2024bnem] estimate the marginal score (or flow) directly and eliminate inner-loop simulation; PINN-based approaches such as NETS [albergo2024nets] similarly train samplers without rollouts. While these advances improve efficiency in continuous spaces, discrete or jump process samplers remain underexplored. Recently, LEAPS [holderrieth2025leaps] introduced a proactive importance sampling scheme for discrete state spaces using locally equivariant networks. However, it requires many energy evaluations at inference since it uses escorted transport.
Generative flow networks (GFNs). GFNs [bengio2021flow, lahlou2023theory] formulate sampling as a sequential decision process: a forward policy constructs an object step by step, and a backward policy enforces consistency so that trajectories terminate with probability proportional to a target reward (or unnormalized density). Training objectives like trajectory balance [malkin2022trajectory] and detailed balance [bengio2023gflownet] enable off-policy updates. Furthermore, falet2023delta propose the sampler for a sparse graphical model with a local objective similar to our forward-looking parametrization. Though diffusion sampler is a special case of a continuous-state continuous-time GFNs [zhang2022unifying, berner2025discrete], it remains an interesting question that training a CTMP on a general state space, as ours, can be unified under the GFNs framework.
Relation to iDEM and BNEM. In EGM, choosing $q_{1|t}(x_{1}|x)\propto p_{t|1}(x|x_{1})$ yields simple weights $\tilde{w}(x,x_{1})=\tilde{p}_{1}(x_{1})=\exp(-\mathcal{E}(x_{1}))$. In the VE path on continuous space where target score identity [de2024target] is available, the integrand can be replaced with the target score instead of the conditional score. This yields the same estimator used in iDEM [akhound2024iterated] and NEM [ouyang2024bnem]. Yet, it remains an open question whether it’s possible to derive the target score identity on a general state space. Recently, zhang2025target introduce the target score identity on discrete states with neighborhood structure, which might improve the accuracy of our estimator on discrete space when combined.
@@ A.1 Importance sampling for the generator estimation
This section provides a detailed derivation of the importance sampling estimator for the generator and its parametrization presented in Equation˜9.
Existence of densities. Consider the measurable space $(S,\Sigma)$ with a $\Sigma$-measurable reference measure $\nu$ (e.g., counting measure if $S$ is discrete and Lebesgue measure if $S=\mathbb{R}^{d}$), as introduced in the Section˜2. Since we focus on a sampling problem, the target distribution $p_{1}$ admits a $\nu$-density $\frac{dp_{1}}{d\nu}:S\to\mathbb{R}_{\geq 0}$. Assume that the joint probability measure $p_{t,1}$ is absolutely continuous with respect to the product measure $\nu\otimes\nu$. Then, all related probability measures $p_{t},p_{t|1},p_{1|t}$ admits corresponding $\nu$-densities, expressed as:
@@
SNIS estimation of the generator. We introduce a proposal distribution $q_{1|t}:\Sigma\times S\to\mathbb{R}_{\geq 0}$, satisfying absolute continuity conditions: $p_{1|t}(\cdot|x)\ll q_{1|t}(\cdot|x)$, $q_{1|t}(\cdot|x)\ll\nu$ and $\nu\ll q_{1|t}(\cdot|x)$ for all $x\in S$. Using the marginalization trick Equation˜3 and the Radon-Nikodym theorem [lang2012real, Chapter 7], we have:
@@ A.2 Proof of Theorem˜1
Theorem 2 (Restatement of Theorem˜1).
@@
The KFE is satisfied for the conditional probability path pt|rp{t|r} by definition:
@@
where the marginal consistency assumption Equation˜13 is applied in the third equality. This concludes the proof. ∎
@@ A.3 Derivation of bootstrapping estimator for generator estimation
We derive a bootstrapping estimator for the marginal generator and its parametrization proposed in Equation˜21, based on the marginalization trick in Equation˜54.
@@
A.4 Analysis on bias and variance of the IS and bootstrapping estimator
In this section, we investigate the bias and variance of the IS and bootstrapping estimator. The goal of the analysis is to support the bootstrapping estimation with a theoretical argument and demonstrate that it provides a lower variance estimator compared to the IS method. To do so, we first derive asymptotic bounds on bias and variance for IS estimation. Then, we derive the bounds for bootstrapping estimation, assuming that the intermediate energy is well-trained. Under this assumption, we show that bootstrapping reduces estimation variance compared to IS, demonstrating that bootstrapping trades off the estimation variance with the model learning bias.
Bias and variance of IS estimation.
We show that the SNIS estimator’s error decays as $O(1/\sqrt{K})$ and its bais and variance decays as $O(1/K)$, where $K$ is the sample size.
Proposition 1.
Consider an unnormalized density $\tilde{p}_{1}(x_{1})$ and a conditional generator $F_{t|1}^{x_{1}}(x_{t})$ evaluated on a sample $x_{1}\sim q_{1|t}$. Suppose $\tilde{p}_{1}(x_{1})$ and $\Big\lVert\tilde{p}_{1}(x_{1})F_{t|1}^{x_{1}}(x_{t})\Big\rVert$ are sub-Gaussian. Then there exists a constant $c(x_{t})$ such that with probability at least $1-\delta$,
$\left\lVert\hat{F}_{t}(x_{t})-F_{t}(x_{t})\right\rVert\leq c(x_{t})\sqrt{\frac{\log(2/\delta)}{K}},$
where $\hat{F}_{t}(x_{t})$ denotes the SNIS estimator from Equation˜9 using $K$ samples $x_{1}^{(1)},\ldots,x_{1}^{(K)}\sim q_{1|t}:$
$\hat{F}_{t}(x_{t})=\frac{\sum_{i=1}^{K}\tilde{p}_{1}(x_{1}^{(i)})F_{t|1}^{x_{1}^{(i)}}(x_{t})}{\sum_{i=1}^{K}\tilde{p}_{1}(x_{1}^{(i)})}.$
Moreover, its bias and variance are expressed as:
$\text{Bias}[\hat{F}_{t}]=\frac{1}{Kp_{t}(x_{t})^{2}}\left(-\text{Cov}\left[\tilde{p}_{1}F_{t|1},\tilde{p}_{1}\right]+F_{t}(x_{t})\text{Var}[\tilde{p}_{1}]\right)+O\left(\frac{1}{K^{2}}\right),$
$\text{Var}[\hat{F}_{t}]=\frac{4\text{Var}[\tilde{p}_{1}(x_{1})]}{p_{t}^{2}(x_{t})K}(1+\lVert F_{t}(x_{t})\rVert)^{2}.$
Proof.
Let $\hat{A}$ and $\hat{B}$ denote the numerator and the denominator of $\hat{F}_{t}$, respectively, i.e.,
$\hat{A}=\frac{1}{K}\sum_{i=1}^{K}\tilde{p}_{1}(x_{1}^{(i)})F_{t|1}^{x_{1}^{(i)}}(x_{t}),\qquad\hat{B}=\sum_{i=1}^{K}\tilde{p}_{1}(x_{1}^{(i)}).$ (89)
We also let $\mathbb{E}[\hat{A}]=A=\tilde{p}_{t}(x_{t})F_{t}(x_{t})$ and $\mathbb{E}[\hat{B}]=B=\tilde{p}_{t}(x_{t})$. By Hoeffding’s inequality for sub-Gaussian random variables, there exists a constant $C$ such that
$\left\lVert\hat{A}-A\right\rVert$ $\leq C\sqrt{\frac{\log(2/\delta)}{K}},\qquad\left|\hat{B}-B\right|\leq C\sqrt{\frac{\log(2/\delta)}{K}},$ (90)
with $1-\delta$ probability. (Here, $C=\sqrt{\text{Var}[\tilde{p}_{1}(x_{1})]}$ is a possible choice.)
Since we have bounds for both numerator and denominator of $\hat{F}_{t}$, we can also bound the error of $\hat{F}_{t}$ to $F_{t}$. The result is as follows:
$\left\lVert\hat{F}_{t}(x_{t})-F_{t}(x_{t})\right\rVert$ $=\left\lVert\frac{\hat{A}}{\hat{B}}-\frac{A}{B}\right\rVert=\left\lVert\frac{\hat{A}B-A\hat{B}}{\hat{B}B}\right\rVert$ (91)
$\leq\frac{\lVert A\rVert\left|\hat{B}-B\right|+B\left\lVert\hat{A}-A\right\rVert}{\hat{B}B}$ (92)
$\leq\frac{1}{\hat{B}B}Cp_{t}(x_{t})(1+\lVert F_{t}(x_{t})\rVert)\sqrt{\frac{\log(2/\delta)}{K}}$ (93)
$\leq\frac{2C}{p_{t}(x_{t})}(1+\lVert F_{t}(x_{t})\rVert)\sqrt{\frac{\log(2/\delta)}{K}}=c(x_{t})\sqrt{\frac{\log(2/\delta)}{K}},$ (94)
where we assume sufficiently large $K$ such that $\hat{B}\geq\frac{1}{2}B$.
Now, for sufficiently large $K$, the Taylor expansion of $\hat{F}_{t}$ is expressed as:
$\hat{F}_{t}=\frac{A}{B}+\frac{1}{B}(\hat{A}-A)-\frac{A}{B^{2}}(\hat{B}-B)-\frac{1}{B^{2}}(\hat{A}-A)(\hat{B}-B)+\frac{A}{B^{3}}(\hat{B}-B)^{2}+O\left(\frac{1}{K^{2}}\right).$
To derive the final equation for the bias term, one can express the bias term as follows:
$\text{Bias}[\hat{F}_{t}]=\mathbb{E}[\hat{F}_{t}]-\frac{A}{B}=-\frac{1}{B^{2}}\text{Cov}[\hat{A},\hat{B}]+\frac{A}{B^{3}}\text{Var}[\hat{B}$].
Since $\text{Cov}[\hat{A},\hat{B}]=\text{Cov}\left[\tilde{p}_{1}F_{t|1},\tilde{p}_{1}\right]/K$ and $\text{Var}[\hat{B}]=\text{Var}[\tilde{p}_{1}]/K$, one obtains the conclusion for $\text{Bias}[\hat{F}_{t}]$.
To derive the final equation for the variance term, one can combine the sub-Gaussianity of $\hat{F}_{t}$ with the error bound on $\left\lVert\hat{F}_{t}(x_{t})-F_{t}(x_{t})\right\rVert$. ∎
Bias and variance of bootstrapping estimation.
Now, we derive analogous asymptotic bounds for the bias and variance of the bootstrapping estimator. We assume a fully trained intermediate energy model, which allows us to isolate the impact of estimation bias on the intermediate energy while disregarding any neural network learning bias.
Proposition 2.
Consider a fully trained surrogate energy model $\mathcal{E}_{r}^{\phi}(x_{r})=\mathbb{E}[\hat{\mathcal{E}}_{r}(x_{r})]$, where $\hat{\mathcal{E}}_{r}(x_{r})$ is a biased energy estimator from Equation˜20. Let $\hat{F}_{t}(x_{t};\mathcal{E}_{r})$ be the bootstrapping estimator using $\mathcal{E}_{r}$ as the intermediate energy function:
$\hat{F}_{t}(x_{t};\mathcal{E}_{r})=\frac{\sum_{i}\exp(-\mathcal{E}_{r}(x_{r}^{(i)}))F_{t|r}^{x_{r}^{(i)}}(x_{t})}{\sum_{i}\exp(-\mathcal{E}_{r}(x_{r}^{(i)}))}.$ (95)
Then, the bias and the variance bound of the bootstrapping estimator with surrogate model $\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})$ are given by:
$\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})]=\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]+O\left(\frac{1}{K^{2}}\right),$
$\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})]=\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]=\frac{\text{Var}[\tilde{p}_{r}(x_{r})]}{\text{Var}[\tilde{p}_{1}(x_{1})]}\text{Var}[\hat{F}_{t}(x_{t})].$
where $\hat{F}_{t}(x_{t})$ is the estimator without bootstrapping. Since $\text{Var}[\tilde{p}_{r}(x_{r})]<\text{Var}[\tilde{p}_{1}(x_{1})]$, the variance of bootstrapping estimator is smaller than the $\hat{F}_{t}(x_{t})$.
Proof.
The bias of the energy estimator $\hat{\mathcal{E}}_{r}$ is given as follows (refer to Corollary 3.2 of [ouyang2024bnem]):
$\text{Bias}[\hat{\mathcal{E}_{r}}(x_{r})]=\frac{\text{Var}[\tilde{p}_{1}(x_{1})]}{2p_{r}^{2}(x_{r})K}+O\left(\frac{1}{K^{2}}\right)$ (96)
Thus, we can approximate the surrogate energy as,
$\mathcal{E}_{r}^{\phi}(x_{r})=\mathbb{E}[\hat{\mathcal{E}}_{r}(x_{r})]=\mathcal{E}_{r}(x_{r})+\underbrace{\frac{\text{Var}[\tilde{p}_{1}(x_{1})]}{2p_{r}^{2}(x_{r})K}}_{:=b(x_{r})}.$ (97)
Plugging the above equation into the $\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})$, we get,
$\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})=\frac{\sum_{i}\exp(-\mathcal{E}^{\phi}_{r}(x_{r}^{(i)}))F_{t|r}^{x_{r}^{(i)}}(x)}{\sum_{i}\exp(-\mathcal{E}^{\phi}_{r}(x_{r}^{(i)}))}=\frac{\sum_{i}\exp(-\mathcal{E}_{r}(x_{r}^{(i)})-b(x_{r}^{(i)}))F_{t|r}^{x_{r}^{(i)}}(x)}{\sum_{i}\exp(-\mathcal{E}_{r}(x_{r}^{(i)})-b(x_{r}^{(i)}))}.$ (98)
Here, $b(x_{r}^{(i)})$ is close to 0 and concentrated to:
$m_{b}=b(x_{t})=\frac{\text{Var}[\tilde{p}_{1}(x_{1})]}{2p_{r}^{2}(x_{t})K},$ (99)
when $t$ is close to $r$ and $K$ is sufficiently large. To keep notation concise, let $w_{i}$ be the self-normalized importance weight with true energy $\mathcal{E}_{r}(x_{r})$ and $F_{t|r}^{(i)}$ be conditional generator with the sample $x_{r}^{(i)}$:
$w_{i}=\frac{\exp(-\mathcal{E}_{r}(x_{r}^{(i)}))}{\sum_{j}\exp(-\mathcal{E}_{r}(x_{r}^{(j)}))},\quad F_{t|r}^{(i)}=F_{t|r}^{x_{r}^{(i)}}(x).$ (100)
Then, by applying first-order Taylor expansion to the $\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})$ and approximation $b(x_{r}^{(i)})\approx m_{b}$, we obtain:
$\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})$ $=\frac{\sum_{i}w_{i}\exp(-b(x_{r}^{(i)}))F_{t|r}^{(i)}}{\sum_{i}w_{i}\exp(-b(x_{r}^{(i)}))}$ (101)
$\approx\frac{\sum_{i}w_{i}(1-b(x_{r}^{(i)}))F_{t|r}^{(i)}}{\sum_{i}w_{i}(1-b(x_{r}^{(i)}))}$ (102)
$=\frac{\sum_{i}w_{i}F_{t|r}^{(i)}-\sum_{i}w_{i}b(x_{r}^{(i)})F_{t|r}^{(i)}}{1-\sum_{i}w_{i}b(x_{r}^{(i)})}$ (103)
$\approx\left(\sum_{i}w_{i}F_{t|r}^{(i)}-\sum_{i}w_{i}b(x_{r}^{(i)})F_{t|r}^{(i)}\right)\left(1+\sum_{i}w_{i}b(x_{r}^{(i)})\right)$ (104)
$\approx\left(\sum_{i}w_{i}F_{t|r}^{(i)}-m_{b}\sum_{i}w_{i}F_{t|r}^{(i)}\right)(1+m_{b})$ (105)
$\approx(1-m_{b}^{2})\frac{\sum_{i}\exp(-\mathcal{E}_{r}(x_{r}^{(i)}))F_{t|r}^{x_{r}^{(i)}}(x)}{\sum_{i}\exp(-\mathcal{E}_{r}(x_{r}^{(i)})))}$ (106)
$=(1-m_{b}^{2})\hat{F}_{t}(x_{t};\mathcal{E}_{r})$ (107)
Therefore, the bias of the bootstrapping estimator $\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})$ is:
$\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})]$ $=\mathbb{E}[\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})]-F_{t}(x_{t})$ (108)
$=(1-m_{b}^{2})\mathbb{E}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]-F_{t}(x_{t})$ (109)
$=(1-m_{b}^{2})(F_{t}+\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})])-F_{t}(x_{t})$ (110)
$=(1-m_{b}^{2})\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]-m_{b}^{2}F_{t}(x_{t})$ (111)
$=\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]-\frac{v_{1r}^{2}(x_{t})}{4p_{r}^{4}(x_{t})K^{2}}\left(F_{t}(x_{t})+\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]\right)$ (112)
$=\text{Bias}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]+O\left(\frac{1}{K^{2}}\right)$ (113)
Similarly, the variance of the bootstrapping estimator is:
$\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})]=(1-m_{b}^{2})^{2}\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]\approx\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]$ (114)
since $m_{b}<1$ for sufficiently large $K$.
The variance of the IS estimator $\hat{F}_{t}(x_{t})$ without bootstrapping is given by:
$\text{Var}[\hat{F}_{t}(x_{t})]=\frac{4\text{Var}[\tilde{p}_{1}(x_{r})]}{p_{t}^{2}(x_{t})K}(1+\lVert F_{t}(x_{t})\rVert)^{2},$ (115)
Also, the variance of the bootstrapping estimator $\hat{F}_{t}(x_{t};\mathcal{E}_{r})$ with true energy is given by:
$\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]=\frac{4\text{Var}[\tilde{p}_{r}(x_{r})]}{p_{t}^{2}(x_{t})K}(1+\lVert F_{t}(x_{t})\rVert)^{2},$ (116)
Consequently,
$\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r}^{\phi})]\approx\text{Var}[\hat{F}_{t}(x_{t};\mathcal{E}_{r})]=\frac{\text{Var}[\tilde{p}_{r}(x_{r})]}{\text{Var}[\tilde{p}_{1}(x_{1})]}\text{Var}[\hat{F}_{t}(x_{t})].$ (117)
Because $\text{Var}[\tilde{p}_{r}(x_{r})]\ll\text{Var}[\tilde{p}_{1}(x_{1})]$, the variance of the bootstrapping estimator is smaller than the $\hat{F}_{t}(x_{t})$. Consequently, with an fully trained energy model, we reduce the variance of the SNIS generator estimation. ∎